OpenStudy (anonymous):

Solve x2 + 6x + 7 = 0. a) x = −1 and x = −5 b) 3 plus or minus square root of 2 c)-3 plus or minus square root of 2 d)-3 plus or minus square root of 2 all over 2

OpenStudy (anonymous):

is that x^2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

use the quadratic formula \[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] A= 1 B = 6 C= 7

OpenStudy (anonymous):

err C

OpenStudy (anonymous):

ya it's -6/2 = -3 and sqrt(8) / 2 = sqrt(2)

OpenStudy (anonymous):

so -3 plus minus sqrt 2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you were right, you just didn't apply the /2

OpenStudy (anonymous):

Np.

OpenStudy (anonymous):

how would i do −16t2 + 64t + 4?

OpenStudy (anonymous):

would it be 2 plus or minus square root of 17 all over 2?

OpenStudy (anonymous):

sorry just saw this

OpenStudy (anonymous):

\[2 \pm \frac{ \sqrt{17} }{ 2 }\]

OpenStudy (anonymous):

since it's 64/32 = 2 it becomes only sqrt(17)/2

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

Np